A coin with P[H]=0.3 is flipped 3 times. Define N to be the total number of heads. Find the CDF FN(n), the cumulative distribution function for N. (please provide an explanation for the answer)
Answer #1

here F(n)= 0   if -\infty < n<1

F(n)=3C0(0.3)0(0.7)3 =0.343 if 0 <=n <1

F(n)=0.343+3C1(0.3)1(0.7)2 =0.343+0.441 =0.784     if 1 <=n<2

F(n)=0.784+3C1(0.3)1(0.7)2 =0.973   if 2<=n <3

F(n)=1 if 3<=n <\infty